OpenStudy (anonymous):

can someone please help me with simplify v^-3(v^-9) write only with positive exponents please help

OpenStudy (anonymous):

I would multiply these powers together first. Can you do that?

OpenStudy (anonymous):

like 3(9)?

OpenStudy (anonymous):

When multiplying powers with the same base, keep the base the same and add the exponents

OpenStudy (anonymous):

so v^12?

OpenStudy (anonymous):

The exponents are -3 and -12. What is their sum?

OpenStudy (anonymous):

21? or -15 but the exponents in the problem are -3 and -9

OpenStudy (anonymous):

Opps sorry. -3 + -9 =?

OpenStudy (anonymous):

-12?

OpenStudy (mathstudent55):

yes

OpenStudy (mathstudent55):

Example: \(\large x^{-5} \times x^{-7} = x^{-5 + (-7)} = x^{-12} \)

OpenStudy (anonymous):

so v^-12? but it wanted a positive exponent

OpenStudy (anonymous):

So now you v^-12. In order to change the sign of the exponent, take the reciprocal of the power. For example, \[x^{-2} = \frac{ 1 }{ x^2 }\]See where to go from here?

OpenStudy (anonymous):

so 1/v^12?

OpenStudy (anonymous):

Correct! Well done

OpenStudy (mathstudent55):

@RaeBeach9296 Correct! Good job! @ospreytriple Good explanations!

OpenStudy (anonymous):

Thanks and good luck with your math studies.

OpenStudy (anonymous):

thank you :-)